OpenStudy (anonymous):

given a graph how do you identify intervals on which f''(x) < 0

OpenStudy (anonymous):

this is a second derivative.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

im looking at a graph however, i dont know how to calculate the intervals on which f''(x) <0

OpenStudy (anonymous):

is this an ad?

OpenStudy (anonymous):

from sammi i think so. i cant afford to pay for help

OpenStudy (anonymous):

im just trying to get help on my homework

OpenStudy (anonymous):

we students may suffer from the budget tension.

OpenStudy (anonymous):

yeah sammi but it cost

OpenStudy (anonymous):

you can earn money by answering questions there too!

OpenStudy (anonymous):

accordingly there are people who pay the money.

OpenStudy (anonymous):

anyway ken could you help me figure this problem out?

OpenStudy (anonymous):

you can do the first derivative , and then do the second.

OpenStudy (anonymous):

for example, suppose f(x)=x^3+6x^2+4x+3 then f'(x)=3x^2+12x+4 f''(x)=6x+12

OpenStudy (anonymous):

how to i find the first der. from the graph

OpenStudy (anonymous):

If you have a graph of f, you look at the places on the graph where the concavity is negative (curve opens downward). This will be the places where the second derivative is negative.

OpenStudy (anonymous):

back up a little bit, do you know the basic derivatives?

OpenStudy (anonymous):

polpak is right, but it is just perceiving by eyes, not calculating

OpenStudy (anonymous):

if i have numbers i can find der., but there is only letters

OpenStudy (anonymous):

to see where the first derivative is negative you look to see where the slope of the tangent to the graph points downward as x moves positively..

OpenStudy (anonymous):

it's the same.

OpenStudy (anonymous):

im not understanding

OpenStudy (anonymous):

f'(x) stands for the slope of the tangent line at (x,f(x)). f''(x) stands for the change of the slope.

OpenStudy (anonymous):

I presume you have some graph of a function, and you need to identify where the second derivative of the function is less than 0

OpenStudy (anonymous):

i do have a graph so where it is sloping below 0

OpenStudy (anonymous):

Is the graph of f, or of f'' ?

OpenStudy (anonymous):

suppose f(x)=ax^3+bx^2+cx+d then f'(x)=3ax^2+2bx+c then f''(x)=6ax+2b

OpenStudy (anonymous):

f'(x)

OpenStudy (anonymous):

If it's the graph of f' then you simply need to see where the tangent lines have negative slope to find where f'' is less than 0.

OpenStudy (anonymous):

i don get ya.(><)

OpenStudy (anonymous):

Hey anyone wants to make money answering questions or do you need homework help? Try www.aceyourcollegeclasses.com